# Microlensing of non-singular black holes at finite size: a ray tracing approach

### v1

@@ Abstract
We study the gravitational microlensing of various static and spherically symmetric non-singular black holes (and horizonless, non-singular compact objects of similar size). For pointlike sources we extend the parametrized post-Newtonian lensing framework to fourth order, whereas for extended sources we develop a ray tracing approach via a simple radiative transfer model. Modelling non-relativistic proper motion of the lens in front of a background star we record the apparent brightness as a function of time, resulting in a photometric lightcurve. Taking the star radius to smaller values, our numerical results approach the theoretical predictions for point-like sources. Compared to the Schwarzschild metric in an otherwise unmodified lensing geometry, we find that non-singular black hole models (and their horizonless, non-singular counterparts) at finite size tend to feature larger magnifications in microlensing lightcurves, contrary to the point-source prediction.
@@ I Introduction
For a light ray trajectory in a black hole spacetime, the majority of gravitational deflection occurs in the direct vicinity of the gravitational radius [7, 8, 9, 10, 11, 12, 13]. For this reason, the deflection of light is rather sensitive to horizon-scale modifications and is under active investigation (see, e.g., [14, 15, 16, 17, 18, 19]) via presently available strong-lensing data stemming from the vicinity of supermassive black holes. However, such astrophysically active regions impede the identification of signatures of new physics [20, 21, 22, 23].
@@
Because microlensing signatures stem from isolated compact gravitational objects, photometric lightcurves are a comparatively straightforward avenue to model deviations from general relativity. A general perturbative formalism to describe lensing in static, spherically symmetric spacetimes has been developed by Keeton and Petters [24], and while strong lensing and black hole shadows have received considerable attention in the context of non-singular black hole models [14, 15, 16, 17, 18, 19], to the best of our knowledge microlensing and photometric lightcurves of non-singular compact objects have not yet been described in detail (see, however, Refs. [25, 26]). In this paper, we will follow a dual approach:
@@
This paper is organized as follows. In Sec. II we will briefly discuss general features of non-singular black hole models, and then describe the models considered in this paper in more detail. In Sec. III we derive the weak gravitational lensing angle and then extend the PPN formulation to fourth order, arriving at an analytical perturbative formula for the microlensing lightcurve applicable to pointlike sources. We then describe a numerical ray tracer in Sec. IV, and in Sec. V we test this ray tracer via a simple accretion disk model around a non-singular black hole in the strong-lensing regime. We turn towards numerical microlensing in Sec. VI by modeling the motion of a rogue black hole that intersects the line of sight between the observer and a distant but extended background star light source. The ray traced intensities for each pixel are then integrated to yield the lightcurve for each scenario, allowing us to quantify the deviations between the Schwarzschild lightcurve prediction and various non-singular black hole models. We also extract the lightcurves of compact horizonless objects, which are obtained from increasing the regularization parameter $\ell$ beyond a critical value. By reducing the size of the light source, we show that the numerical results converge to the PPN expressions valid only for pointlike sources. We conclude our findings in Sec. VII and demonstr. In appendix A we list complete PPN formulas to fourth order, and the numerical information for the lightcurve simulations as well as additional ray traced images are summarized in appendix B.
@@ III.2 Fourth-order PPN expressions
$r_{\text{E}}$ $=d_{\text{os}}\tan\vartheta_{E}=d_{\text{os}}\tan\sqrt{\frac{4GM\,d_{\text{ls}}}{d_{\text{ol}}d_{\text{os}}}}\,.$ (18)
@@
$C(r)=\tilde{r}(r)\,,\quad\frac{\mbox{d}r}{B(r)}=\tilde{B}(\tilde{r})\mbox{d}\tilde{r}\,$, (22)
@@
and find [^2] We note that Keeton and Petters [24] report on an additional term $16b_{3}/3$ in their expression for $A_{3}$ which we cannot reproduce.
@@
$A_{3}$ $=\frac{70}{3}a_{1}^{3}-20a_{1}a_{2}+4a_{3}+10a_{1}^{2}b_{1}-4a_{2}b_{1}$
$\hskip 11.0pt-2a_{1}b_{1}^{2}+\frac{2}{3}b_{1}^{3}+8a_{1}b_{2}-\frac{8}{3}b_{1}b_{2}\,$,
$A_{4}$ $=\bigg(30a_{1}^{4}-36a_{1}^{2}a_{2}+\frac{9}{2}a_{2}^{2}+9a_{1}a_{3}-\frac{3}{2}a_{4}$
@@
$\hskip 11.0pt+\frac{3}{4}a_{1}b_{3}^{2}-\frac{15}{64}b_{1}^{4}+9a_{1}^{2}b_{2}-3a_{2}b_{2}-3a_{1}b_{1}b_{2}$
$\hskip 11.0pt+\frac{9}{8}b_{1}^{2}b_{2}-\frac{3}{4}b_{2}^{2}-\frac{3}{2}b_{3}+3b_{4}\bigg)\pi\,.$
@@
Noting that $b=d_{\text{ol}}\sin\theta$, we can insert the above relations into the lensing equation (15) for the general form of the deflection angle (29) to eliminate the expansion factors $GM/b$, arriving at an expansion in $\epsilon/\theta_{0}$ instead:
$\alpha$ $=A_{1}\frac{\epsilon}{\theta_{0}}+(A_{2}-A_{1})\,\theta_{1}\left(\frac{\epsilon}{\theta_{0}}\right)^{2}$ (34)
@@
with $d\equiv d_{\text{ol}}/d_{\text{os}}$. Inserting the above results back into the lensing equation (15), we need to recover a null result for consistency, which we can then use to express the image positions $\theta_{i}$ in terms of the constants $A_{i}$. We find
@@ III.3 OGLE-2011-BLG-0462/MOA-2011-BLG-191
Recently, a stellar-mass rogue black hole has been detected via microlensing [41, 42, 43]; the estimated parameters of this event are
@@ III.4 PPN lightcurves
The results are qualitatively similar between different non-singular models which is why we focus on the Minkowski core example. Since $a_{1}=b_{1}=1$ for all models considered, $\mu_{\text{tot,0}}$ agrees between all models since it only depends on $A_{1}=2(a_{1}+b_{1})$. As expected, the next-to-leading-order contributions are heavily suppressed, but nevertheless distinct from the Schwarzschild case. In the point-source approximation, the net effect is a decrease in magnification.
Figure 3: The parametrized post-Newtonian lightcurves (zeroth-order contributions: solid line; second-order contributions: dashed lines; third-order contributions: dotted lines) for the parameters described in the text. Note that the zeroth-order curve is rescaled by a factor of $1/20$. Higher order PPN corrections reduce the magnification, in line with the intuition from the fact that non-singular black holes generally have smaller horizons than their equal-mass Schwarzschild counterparts.
@@ IV Ray tracing for static, spherically symmetric and asymptotically flat metrics
With the perturbative machinery developed, let us now discuss the numerical approach. In static, spherically symmetric spacetimes (3)–(4) we may utilize the existence of a Killing vectors to introduce constants of motion. In asymptotically flat spacetimes we may utilize Schwarzschild coordinates $\{t,r,\theta,\varphi\}$, wherein the Killing vectors are given by
@@
It is hence clear that this system is integrable, but for numerical evaluation not all equivalent formulations are equally efficient. Since we are only interested in the spatial aspects of the problem, we use the affine freedom of null geodesics and rescale their affine parameter such that $E=1$ and subsequently eliminate the equation of motion for $t(\lambda)$. Working with Mathematica, we find that the utilization of $L_{1}$ to render the equation of motion for $\varphi(t)$ first order leads to significant speedup, whereas implementing $L_{2}$ and $L_{3}$ (to obtain first-order equations for $r(\lambda)$ and $\theta(\lambda)$) actually slows the code down, most likely due to the non-linear nature of the radial equation as well as the more involved appearance of trigonometric functions in the polar equation. Moreover, the possible existence of multiple turning points in $r$ make it inconvenient to rephrase the radial differential equation in terms of an effective potential.
@@ V.2 Accretion disks
Similar to the previously discussed case we will again utilize a simple, phenomenologically motivated model for an accretion disk as found, for example, in Ref. [47], taking the form
@@ VII Summary and conclusions
This paper describes, for the first time, the microlensing of non-singular black holes (and their compact, horizonless counterparts) for finite-size sources. After extending the work by Keeton and Petters in the parametrized post-Newtonian framework to fourth order, we demonstrated that our numerical approach reproduces the theoretical predictions as the size of the source shrinks, as expected. Taking this as a validation of our numerical setup, we then explored the resulting finite-source microlensing lightcurves for various non-singular black hole models in the near-zone and far-zone regime. We find throughout that the existence of a regulator length scale $\ell>0$ tends to increase the magnification recorded in finite-source lightcurves, contrary to the point-source prediction.

### v2

@@ Abstract
We study the gravitational microlensing of various static and spherically symmetric non-singular black holes (and horizonless, non-singular compact objects of similar size). For pointlike sources we extend the parametrized post-Newtonian lensing framework to fourth order, whereas for extended sources we develop a ray tracing approach via a simple radiative transfer model. Modelling non-relativistic proper motion of the lens in front of a background star we record the apparent brightness as a function of time, resulting in a photometric lightcurve. Taking the star radius to smaller values, our numerical results approach the theoretical predictions for point-like sources. Compared to the Schwarzschild metric in an otherwise unmodified lensing geometry, we find that non-singular black hole models (and their horizonless, non-singular counterparts) both at finite size and for point lenses tend to feature larger magnifications in microlensing lightcurves.
@@ I Introduction
For a light ray trajectory in a black hole spacetime, the majority of gravitational deflection occurs in the direct vicinity of the gravitational radius [7, 8, 9, 10, 11, 12, 13]. For this reason, the deflection of light is rather sensitive to horizon-scale modifications and is under active investigation (see, e.g., [14, 15, 16, 17, 18, 19, 20, 21]) via presently available strong-lensing data stemming from the vicinity of supermassive black holes. However, such astrophysically active regions impede the identification of signatures of new physics [22, 23, 24, 25].
@@
Because microlensing signatures stem from isolated compact gravitational objects, photometric lightcurves are a comparatively straightforward avenue to model deviations from general relativity. A general perturbative formalism to describe lensing in static, spherically symmetric spacetimes has been developed by Keeton and Petters [26], and while strong lensing and black hole shadows have received considerable attention in the context of non-singular black hole models [14, 15, 16, 17, 18, 19, 20, 21], to the best of our knowledge microlensing and photometric lightcurves of non-singular compact objects have not yet been described in detail (see, however, Refs. [27, 28, 29, 30, 31, 32, 33]). In this paper, we will follow a dual approach:
@@
This paper is organized as follows. In Sec. II we will briefly discuss general features of non-singular black hole models, and then describe the models considered in this paper in more detail. In Sec. III we derive the weak gravitational lensing angle and then extend the PPN formulation to fourth order, arriving at an analytical perturbative formula for the microlensing lightcurve applicable to pointlike sources. We then describe a numerical ray tracer in Sec. IV, and in Sec. V we test this ray tracer via a simple accretion disk model around a non-singular black hole in the strong-lensing regime. We turn towards numerical microlensing in Sec. VI by modeling the motion of a rogue black hole that intersects the line of sight between the observer and a distant but extended background star light source. The ray traced intensities for each pixel are then integrated to yield the lightcurve for each scenario, allowing us to quantify the deviations between the Schwarzschild lightcurve prediction and various non-singular black hole models. We also extract the lightcurves of compact horizonless objects, which are obtained from increasing the regularization parameter $\ell$ beyond a critical value. By reducing the size of the light source, we show that the numerical results converge to the PPN expressions valid only for pointlike sources. We conclude our findings in Sec. VII and in appendix A we list complete PPN formulas to fourth order, and the numerical information for the lightcurve simulations as well as additional ray traced images are summarized in appendix B.
@@ III.2 Fourth-order PPN expressions
$r_{\text{E}}$ $=d_{\text{ol}}\tan\vartheta_{E}=d_{\text{os}}\tan\sqrt{\frac{4GM\,d_{\text{ls}}}{d_{\text{ol}}d_{\text{os}}}}\,.$ (18)
@@
$C(r)=\tilde{r}(r)\,,\quad\frac{\mbox{d}r^{2}}{B(r)}=\tilde{B}(\tilde{r})\mbox{d}\tilde{r}^{2}\,$, (22)
@@
and find
@@
$A_{3}$ $=\frac{70}{3}a_{1}^{3}-20a_{1}a_{2}+4a_{3}+10a_{1}^{2}b_{1}-4a_{2}b_{1}$ (32)
$\hskip 11.0pt-2a_{1}b_{1}^{2}+\frac{2}{3}b_{1}^{3}+8a_{1}b_{2}-\frac{8}{3}b_{1}b_{2}+\frac{16}{3}b_{3}\,$,
$A_{4}$ $=\bigg(30a_{1}^{4}-36a_{1}^{2}a_{2}+\frac{9}{2}a_{2}^{2}+9a_{1}a_{3}-\frac{3}{2}a_{4}$ (33)
@@
$\hskip 11.0pt+\frac{3}{4}a_{1}b_{1}^{3}-\frac{15}{64}b_{1}^{4}+9a_{1}^{2}b_{2}-3a_{2}b_{2}-3a_{1}b_{1}b_{2}$
$\hskip 11.0pt+\frac{9}{8}b_{1}^{2}b_{2}-\frac{3}{4}b_{2}^{2}-\frac{3}{2}b_{1}b_{3}+6a_{1}b_{3}+3b_{4}\bigg)\pi\,.$
@@
Noting that $b=d_{\text{ol}}\sin\vartheta$, we can insert the above relations into the lensing equation (15) for the general form of the deflection angle (29) to eliminate the expansion factors $GM/b$, arriving at an expansion in $\epsilon/\theta_{0}$ instead:
$\alpha$ $=A_{1}\left(\frac{\epsilon}{\theta_{0}}\right)+(A_{2}-A_{1}\theta_{1})\,\left(\frac{\epsilon}{\theta_{0}}\right)^{2}$ (36)
@@
with $d\equiv d_{\text{ls}}/d_{\text{os}}$. Inserting the above results back into the lensing equation (15), we need to recover a null result for consistency, which we can then use to express the image positions $\theta_{i}$ in terms of the constants $A_{i}$. We find
@@
Let us now isolate the leading order impact of the regulator $\ell>0$ on the magnification. We define the dimensionless quantity
$\xi=\frac{\ell}{GM}\,$, (47)
which captures the deviations from general relativity (see Table 2 for the PPN expansion coefficients $a_{i}$ and $b_{i}$ of the non-singular metrics discussed in this work). We then expand the difference between the total magnification (45) for the Schwarzschild metric (denoted “S”) and the Hayward metric, Minkowski core metric, and Simpson–Visser metric (denoted “H,” “M,” and “W,” respectively) in $\xi$ to leading order and find
$\delta\mu_{\text{SH}}$ $\equiv\mu_{\text{tot,S}}(\beta)-\mu_{\text{tot,H}}(\beta)$ (48)
$=-\frac{15\pi\epsilon^{3}\xi^{2}}{16\beta}+\mathcal{O}\left(\xi^{3}\right)\,$,
$\delta\mu_{\text{SM}}$ $\equiv\mu_{\text{tot,S}}(\beta)-\mu_{\text{tot,M}}(\beta)$ (49)
$=\left(-\frac{\epsilon^{2}\xi}{6144\beta}\right)\Bigg[\frac{196608}{(4+\beta^{2})^{3/2}}-\frac{51840\pi^{2}}{(4+\beta^{2})^{5/2}}$
$\hskip 20.0pt-\left(38304+18432d^{2}-6075\pi^{2}\right)\pi\epsilon\Bigg]+\mathcal{O}(\xi^{2})\,$,
$\delta\mu_{\text{SW}}$ $\equiv\mu_{\text{tot,S}}(\beta)-\mu_{\text{tot,W}}(\beta)$ (50)
$=\left(\frac{\epsilon^{2}\xi^{2}}{12288\beta}\right)\Bigg[\frac{32768}{(4+\beta^{2})^{3/2}}-\frac{17280\pi^{2}}{(4+\beta^{2})^{5/2}}$
$\hskip 20.0pt-\left(17568+6144d^{2}-2025\pi^{2}\right)\pi\epsilon\Bigg]+\mathcal{O}(\xi^{3})\,.$
Note that depending on the non-singular metric under consideration, the correction terms enter quadratic order in $\epsilon$ (for the cases M and W) or at cubic order in $\epsilon$ (for the well-known case H). Conversely, their dependence on the dimensionless regulator $\xi$ is linear (for case M) or quadratic (for the cases H and W).
Once the lensing geometry is known, the dimensionless source angle $\beta=\mathcal{B}/\vartheta_{\text{E}}$ can be inserted in order to estimate the quantitative result. As expected, these expressions are suppressed by powers of $\epsilon$ and $\xi$. However, we emphasize that they scale with $1/\beta$, which can amplify them for microlensing scenarios in which the source angle is small. In this case, our formulas (48)–(50) can be used to directly constrain the new physics parameter $\ell$ (given the prior of a non-singular metric). In terms of dimensionful parameters,
$\begin{split}\frac{\epsilon^{p}}{\beta}&=\left[\sqrt{\frac{d_{\text{ol}}d_{\text{os}}}{4GM\,d_{\text{ls}}}}\arctan\left(\frac{GM}{d_{\text{ol}}}\right)\right]^{p}\frac{\vartheta_{\text{E}}}{\mathcal{B}}\\ &\approx\frac{1}{2^{p-1}}\sqrt{\frac{(GM)^{p+1}\,d_{\text{os}}^{p-1}}{d_{\text{ol}}^{p+1}d_{\text{ls}}^{p-1}}}\,\frac{1}{\mathcal{B}}\,,\quad p=2,3\,.\end{split}$ (51)
If the above formulas are applied at the moment of smallest effective source angle, the result is independent of the black hole proper motion parameters, which further simplifies the considerations.
In contrast, other analyses of the perturbative expressions (45), such as the study of the full-width-half-maximum (FWHM), are exacerbated by the non-trivial $\beta$-dependence and are sensitive to the black hole proper motion parameters. For this reason, detailed studies involving astrometric parameters of the lens itself lie beyond the scope of this work.
@@ III.3 OGLE-2011-BLG-0462/MOA-2011-BLG-191
In order to estimate the size of these effects, let us apply the developed formulae to a microlensing event. Recently, a stellar-mass rogue black hole has been detected via microlensing [48, 49, 50]; the estimated parameters of this event are
@@
We extract the minimal angle of $\beta_{T=0.5}\approx 0.05$ and insert this into the expressions (48)–(50), yielding
$\delta\mu_{SH}$ $=-1.9\times 10^{-24}\,\xi^{2}\,\,\mu_{S}\,$, (58)
$\delta\mu_{SM}$ $=-5.4\times 10^{-17}\,\xi\,\mu_{S}\,$, (59)
$\delta\mu_{SW}$ $=-2.0\times 10^{-17}\,\xi^{2}\,\mu_{S}\,$, (60)
where we normalized with respect to the third-order Schwarzschild magnification $\mu_{S}$. The minus signs in the above expressions mean that non-singular metrics feature a larger magnification. For this microlensing event the relative magnification increase is very small, even if $\xi\sim 1$, largely due to the small $\epsilon\sim 8.6\times 10^{-9}$.
@@ III.4 PPN lightcurves
The results are qualitatively similar between different non-singular models which is why we focus on the Minkowski core example. Since $a_{1}=b_{1}=1$ for all models considered, $\mu_{\text{tot,0}}$ agrees between all models since it only depends on $A_{1}=2(a_{1}+b_{1})$. As expected, the next-to-leading-order contributions are heavily suppressed, but nevertheless distinct from the Schwarzschild case. Computing the relative magnification deviations, we find for this numerical setup
$\delta\mu_{SH}$ $=-0.02\,\xi^{2}\,\,\mu_{S}\,$, (63)
$\delta\mu_{SM}$ $=-0.15\,\xi\,\mu_{S}\,$, (64)
$\delta\mu_{SW}$ $=-0.01\,\xi^{2}\,\mu_{S}\,$, (65)
corresponding to percent-level increases in magnification. This again shows that the regulator $\ell$ increases magnification compared to the Schwarzschild case.
Last, we point out that for larger regulators the black hole horizon disappears, resulting in a horizonless geometry. Then, the PPN magnification setup is no longer applicable since the horizon scale is no longer of physical significance. We take this as another argument to develop a numerical setup towards microlensing.
Figure 3: The parametrized post-Newtonian (PPN) lightcurves (zeroth-order contributions: solid line; second-order contributions: dashed lines; third-order contributions: dotted lines) for the three cases of the Schwarzschild metric (S), the Minkowski core metric with a horizon ($M_{-}$) and without a horizon ($M_{+}$), for the parameters described in the text. Note that the zeroth-order curve is rescaled by a factor of $1/20$. Higher order PPN corrections reduce the magnification, but the presence of a regulator reduces that amount, leading to a net increase in magnification.
@@ IV Ray tracing for static, spherically symmetric and asymptotically flat metrics
With the perturbative machinery developed, let us now discuss a numerical approach. In static, spherically symmetric spacetimes (3)–(4) we may utilize the existence of a Killing vectors to introduce constants of motion. In asymptotically flat spacetimes we may utilize Schwarzschild coordinates $\{t,r,\theta,\varphi\}$, wherein the Killing vectors are given by
@@
It is hence clear that this system is integrable, but for numerical evaluation not all equivalent formulations are equally efficient. Since we are only interested in the spatial aspects of the problem, we use the affine freedom of null geodesics and rescale their affine parameter such that $E=1$ and subsequently eliminate the equation of motion for $t(\lambda)$. It is possible to work with first-order differential equations via effective potentials, as done in the pioneering work by Luminet [53] for the imaging of accretion disks and that of Rauch and Blandford [54] in the context of gravitational lensing of Kerr black holes. Working with Mathematica, we find that the utilization of $L_{1}$ to render the equation of motion for $\varphi(\lambda)$ first order leads to significant speedup, whereas implementing $L_{2}$ and $L_{3}$ (to obtain first-order equations for $r(\lambda)$ and $\theta(\lambda)$) actually slows the code down, most likely due to the non-linear nature of the radial equation as well as the more involved appearance of trigonometric functions in the polar equation. Moreover, the possible existence of multiple turning points in $r$ make it inconvenient to rephrase the radial differential equation in terms of an effective potential without previously identifying the location of turning points, which would have to be performed in a distinct way for each metric considered in this work. For this reason we opt to utilize the second-order approach in the radial and polar sector.
@@
In order to accommodate arbitrary, three-dimensional densities for the radiative transport equation we chose not to eliminate $\theta$ entirely. While allowed due to the spherical symmetry of the black hole geometries under consideration in this paper, it would still necessitate the computation of the full three-dimensional trajectory due to the explicit dependence on all three Cartesian coordinates $x$, $y$, and $z$ in the radiation density. The effective polar angle would still be different for each ray-traced pixel (since it is determined by the plane of the black hole, the observer, and the pixel) which would lead to more numerical overhead. In practice, a resolution of $200\times 200$ pixel is sufficient to generate images of sufficient accuracy, and we have verified that images of a larger resolution of $1000\times 1000$ only provide marginal improvement. Since the primary goal of the numerical aspect of this work is to generate ray traced lightcurves, an enhanced angular resolution to capture faint details is not required. Moreover, no large-scale data analysis is performed, and for this reason the full, three-dimensional treatment is retained. Last, this will also streamline the future implementation of axially symmetric metrics with angular momentum in a straightforward manner.
@@ V.2 Accretion disks
Similar to the previously discussed case we will again utilize a simple, phenomenologically motivated model for an accretion disk as found, for example, in Ref. [55, 56], taking the form
@@ VII Summary and conclusions
This paper describes, for the first time, the microlensing of non-singular black holes (and their compact, horizonless counterparts) for finite-size sources. After extending the work by Keeton and Petters in the parametrized post-Newtonian framework to fourth order, we demonstrated that our numerical approach reproduces the theoretical predictions as the size of the source shrinks, as expected. Taking this as a validation of our numerical setup, we then explored the resulting finite-source microlensing lightcurves for various non-singular black hole models in the near-zone and far-zone regime. We find throughout that the existence of a regulator length scale $\ell>0$ tends to increase the magnification recorded in finite-source lightcurves, similar to the point-source predictions.
This is consistent with theoretical expectations: [^2] We thank our Referee for bringing this to our attention. Within the description of gravitational lensing via Fermat’s principle, the image magnification is proportional to the inverse curvature of the local time-delay surface (see e.g. [7] for a pedagogical reference as well as Refs. [57, 58, 59]). Since that curvature is smaller for smoothed-out matter distributions (such as the non-singular metrics considered in the present paper), this would then lead to a larger magnification.